You can afford monthly payments of $200 into an account that pays 3.3% compounded monthly. How long will it be until you have $12,000 to buy a boat?
Answer #1

Future Value = Periodical Payments * \frac{(1+r)^{n}-1}{r}

where r is the rate of return of 1 period i.e 3.3%/12 = 0.275% OR 0.00275

n is the no of compounding periods

12,000= 200* \frac{(1+0.00275)^{n}-1}{0.00275}

(1+0.00275)^{n} = \frac{12000*0.00275}{200} +1

(1+0.00275)^{n} = 1.165

Taking LOG both sides

n * log (1.00275) = log 1.165

n*0.00119 = 0.06633

n = 0.06633/ 0.00119

n = 55.75 years

It will take 55.75 years

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.